OpenStudy (anonymous):

Solve three-fourths b minus one-sixth equals one-half for b. b equals eight-ninths b equals one-half b equals one-fourth b equals four-ninths

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }b-\frac{ 1 }{ 6 }=\frac{ 1 }{ 2 }b\] multyply by 12 \[9b-2=6b\implies 3b=2,b=2/3\]

OpenStudy (anonymous):

@Jonask ty!

OpenStudy (anonymous):

@Jonask what next

OpenStudy (anonymous):

but its none of the options

OpenStudy (anonymous):

@Jonask you made a mistake I think